OpenStudy (anonymous):

Potential difference mean the difference of force acting on the charge between 2 different point?

OpenStudy (anonymous):

 if the reading is v. Can i say that the force, v push the charge forward? is that true?

OpenStudy (stormfire1):

That explanation is close but technically volts = N*m/C = J/C which is to say that voltage is most closely related to work/coulomb.